A nunber is 2/5 times another number. If their sum is 70,find the numbers.
Answer by rfer(16322)

x+2/5x=70
1 2/5x=70
x=70/(1 2/5)
x=70/(7/5)
x=70*5/7
x=50
2/5x=20